OpenStudy (anonymous):

Faces:8 vertices:12 edges:n Use euler's theorem to find the value of n.

OpenStudy (anonymous):

F + V = E + 2 8 + 12 = E + 2 20 = E + 2 E = 18

OpenStudy (anonymous):

thank you so much:)

OpenStudy (anonymous):

You're welcome :)